# Quantum-enhanced sensing of spin-orbit coupling without fine-tuning

### v2

@@ 1 Introduction
The original notion of Spin-Orbit Coupling (SOC) refers to the relativistic quantum mechanical effect, connecting the spin and motional degrees of freedom of electrons in an electric field. SOC plays a crucial role in solid state systems by affecting the energy spectrum in ways that has led to major research areas such as spintronics [1, 2], topological insulators [3, 4], quantum dot arrays [5, 6, 7, 8] and quantum simulation with cold gases under artificial gauge fields [9, 10, 11]. In particular, the Rashba type of SOC [12], which may occur naturally in a solid state material [13] or can be engineered in quantum systems [14, 15, 16], plays a crucial role in a wide range of technological applications. This includes spin field-effect transistors [17], memory and logic devices [18, 19], spin-orbit torque devices [20, 21], spin filters and pumps [22], dissipation-less spin transport [23] and interaction-based quantum transport [24]. In addition, Rashba SOC plays a significant role in emerging quantum technologies from solid state quantum simulators [25, 26, 27, 28] to fault-tolerant quantum computation through formation of Majorana fermions [29] and information encoding in pseudo-spins [30]. Therefore, precise knowledge of the SOC parameters is necessary for studying both equilibrium and non-equilibrium properties of the system. So far, SOC has been measured using various methods such as, electron transport [31, 32], spectroscopy [33, 34], and spin Hall effect [35]. The fundamental and technological importance of SOC and the recent developments in the field of quantum sensing pose a timely question: can the precision, with which the SOC parameters are measured, be enhanced using quantum features?
@@
Quantum sensing has now established itself as a key component of quantum technologies that can surpass the limits of traditional classical sensors [36]. The Cramér-Rao inequality from estimation theory provides a lower bound on the precision of estimating an unknown parameter in terms of inverse of the Fisher information. Generally, as the probe size $L$ is increased, Fisher information grows as $F{\sim}L^{\beta}$. For classical probes, the best achievable scaling is the standard limit where $\beta{=}1$. Meanwhile, exploiting quantum features can improve the precision beyond the capacity of classical probes, with $\beta{>}1$ [37, 36, 38], which is known as quantum-enhanced sensitivity. The concept of quantum-enhanced sensitivity was first introduced in an interferometric setup utilizing Greenberger–Horne–Zeilinger (GHZ) type entangled states [39, 40, 41]. It significantly improves the precision of phase shift measurements to achieve the Heisenberg limit, $\beta{=}2$. Since then, this phenomenon has been verified in various platforms including ion-traps [42], optical systems [43, 44, 45], superconducting qubits [46] and nitrogen-vacancy centers [47]. However, the approach exclusively relies on GHZ-type entangled states, which are challenging to generate and highly susceptible to decoherence and particle loss. Furthermore, this method is limited to a phase generating operation to encode information into the quantum state of the probe. Any disturbance in this unitary process reduces the precision, therefore limits the interferometric sensing method to scale up effectively [48].
@@
Another approach to quantum sensing utilizes many-body probes by leveraging a variety of quantum features. Although interactions have destructive effects in the GHZ-based sensing schemes, they play the primary role in many-body sensors [49]. In particular, quantum criticality has been identified as a resource for achieving quantum enhanced sensitivity. This includes various types of criticalities, such as first-order [50, 51, 52, 53, 54], second-order  [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], Floquet [67, 68], time crystal [69, 70, 71, 72, 73], Stark [74, 75, 76] and quasi-periodic [77] localization, and topological [78, 79, 80] phase transitions. In all these criticalities, a common theme that emerges as the resource for quantum enhancement of the sensing capability is the closing of the energy gap at the phase transition point [49]. Nonetheless, the phase transition occurs only at a particular value of a Hamiltonian parameter which restricts the quantum advantage to a small region around the critical point. As one moves away from the criticality, the energy gap reopens and the quantum enhancement starts to disappear, typically falling back to the standard limit of sensing. This necessitates fine-tuning, making it primarily useful for local sensing where substantial prior knowledge about the parameter of interest is required. Therefore, it is desirable to have a quantum system that feature the gap closing behavior over a wide range of the parameter to be estimated.
@@
2 Overview: parameter estimation
To estimate $d$ number of unknown parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})$, the first step is to encode them in a quantum probe whose density operator $\rho_{\bm{\alpha}}$ depends on $\bm{\alpha}$. To determine these parameters, one needs to perform a measurement, which is represented by a set of Positive Operator-Valued Measure (POVM) operators $\{\Pi_{k}\}$ applied to the probe. The outcomes follow a classical probability distribution, where the probability of the $k$-th outcome is given by $p_{k}(\bm{\alpha})=\mathrm{Tr}[\Pi_{k}\rho_{\bm{\alpha}}]$. These outcomes are fed into an estimator function to infer the parameters. For this measurement setup, the uncertainty in estimating $\bm{\alpha}$ can be quantified by the covariance matrix $[\text{Cov}({\bm{\alpha}})]_{i,j}{=}\braket{\alpha_{i}\alpha_{j}}-\braket{\alpha_{i}}\braket{\alpha_{j}}$ Note that the diagonal elements are the single parameter variances $\sigma^{2}_{\alpha_{i}}$. The lower bound in uncertainty is set by the Craémer-Rao inequality [81, 82, 83, 84]
@@ 2 Overview: parameter estimation
where $M$ is the number of performed measurements and $\mathcal{F}^{C}$ is the $d\times d$ classical Fisher information matrix (CFIM) with elements
@@
This basis-dependent lower bound can be optimized over all possible measurements to give the ultimate obtainable precision in terms of the quantum Cramér-Rao bound [84]
@@
The quantum Fisher information matrix (QFIM) elements can be expressed in terms of the eigen-decomposition of the density operator $\rho_{\bm{\alpha}}{=}\sum_{n,(\lambda_{n}\neq 0)}\lambda_{n}\ket{\lambda_{n}}\bra{\lambda_{n}}$ | in the following way [84],
@@
$+\sum_{n,(\lambda_{n}\neq 0)}4\lambda_{n}\text{Re}(\braket{\partial_{\alpha_{i}}\lambda_{n}}{\partial_{\alpha_{j}}\lambda_{n}})$
@@
$-\sum_{\begin{subarray}{c}n,m\\ (\lambda_{n},\lambda_{m}\neq 0)\end{subarray}}8\frac{\lambda_{n}\lambda_{m}}{\lambda_{n}+\lambda_{m}}\text{Re}(\braket{\partial_{\alpha_{i}}\lambda_{n}}{\lambda_{m}}\braket{\lambda_{m}}{\partial_{\alpha_{j}}\lambda_{n}}).$ (4)
For a pure state $\rho_{\bm{\alpha}}=\ket{\psi_{\bm{\alpha}}}\bra{\psi_{\bm{\alpha}}}$ |, QFIM becomes
$\mathcal{F}^{Q}_{i,j}=4\text{Re}(\braket{\partial_{\alpha_{i}}\psi}{\partial_{\alpha_{j}}\psi})-\braket{\partial_{\alpha_{i}}\psi}{\psi}\braket{\psi}{\partial_{\alpha_{j}}\psi}).$ (5)
@@
To extract a scalar form from the matrix inequalities, a positive weight matrix $W$ can be multiplied on both sides, followed by a trace operation. In particular, for the choice of $W$ being an identity matrix one gets $\sum_{i}\sigma^{2}_{\alpha_{i}}\geq\text{Tr}((\mathcal{F}^{Q})^{-1})/M$. Another useful choice is to take a $W$ matrix with all elements being zero except $W_{i,i}{=}1$. Then the Cramér-Rao inequality becomes $\sigma^{2}_{\alpha_{i}}\geq(\mathcal{F}^{Q})^{-1}_{i,i}/M\geq 1/M\mathcal{F}^{Q}_{i,i}$, in which the last term is associated with the uncertainty in $\alpha_{i}$ when all the other parameters are known. This situation is known as the single parameter estimation problem where the error in estimating one unknown parameter $\alpha$ is lower bounded by $\sigma^{2}_{\alpha}\geq 1/MF^{C}(\alpha)\geq 1/MF^{Q}(\alpha)$. Here, $F^{C}(\alpha)$ and $F^{Q}(\alpha)$ are the single parameter classical Fisher information (CFI) and quantum Fisher information (QFI), and are just the diagonal elements in CFIM and QFIM, respectively. One choice of optimal basis, in which CFI saturates the QFI, is given by the projcetors to the eigenstates of the so called symmetric logarithmic operator (SLD) $\mathcal{L}$, which are implicitly defined for a probe state $\rho$ as $\partial_{\alpha}\rho{=}(\rho\mathcal{L}{+}\mathcal{L}\rho)/2$.
3 Model: 1D quantum wire
@@
The model considered in this work is based on the transport of electrons in a one-dimensional ballistic quantum wire in the presence of Rashba SOC [85, 86]. Low dimensional transport properties of electrons are of extreme importance in condensed matter physics, with applications in spintronics and quantum information processing. SOC is quite prevalent in two-dimensional electron gas systems with strong confinement along the $z$-axis. Further confinement to a wire geometry enhances the effect of SOC [17] and has led to extensive studies on the transports of non-interacting electrons in quasi 1D systems. The effective dynamics can be described by a tight-binding lattice model, which is governed by the Hamiltonian in the following form
@@
[…]
@@
4 Single Parameter Sensing
@@
In low-dimensional transport experiments, accurate knowledge of the SOC parameters is desired. Therefore, we first focus on estimating the SOC parameter in the uniform Rashba field scenario along $y$- and $z$-direction, i.e. $\alpha_{y}{=}\alpha_{z}$. The non-uniform SOC case will be addressed in the next section. The goal is to analyze how well one can estimate $\alpha_{z}$ using the ground state of the system as a probe. To stay close to the experimental systems, we consider finite wires with $L$ lattice sites and study the growth of QFI with system size. As the scaling of QFI is closely connected to the gap closing of the system, we also analyze the energy gap $\Delta$ between first excited state and ground state. The ground state is doubly degenerate, so a Zeeman term is essential to break the degeneracy. This term need to be strong enough to capture the actual energy gap and not the Zeeman splitting. For example, one should choose $B{>}0.005t$ for $L{=}100$.
4.1 Single particle probe

### v3

# Quantum-enhanced sensing of spin-orbit coupling without fine tuning

@@ I Introduction
The original notion of Spin-Orbit Coupling (SOC) refers to the relativistic quantum mechanical effect, connecting the spin and motional degrees of freedom of electrons in an electric field. SOC plays a crucial role in solid state systems by affecting the energy spectrum in ways that has led to major research areas such as spintronics Nagaosa et al. (2010); Hirohata et al. (2020), topological insulators Hasan and Kane (2010); Qi and Zhang (2011), quantum dot arrays Golovach et al. (2004); Bulaev et al. (2008); Tanttu et al. (2019); Iijima and Akera (2020) and quantum simulation with cold gases under artificial gauge fields Dalibard et al. (2011); Goldman et al. (2014); Zhang and Liu (2018). In particular, the Rashba type of SOC Manchon et al. (2015), which may occur naturally in a solid state material Min et al. (2006) or can be engineered in quantum systems Zheng et al. (2015); Governale (2002); Governale and Zülicke (2002), plays a crucial role in a wide range of technological applications. This includes spin field-effect transistors Datta and Das (1990), memory and logic devices Miron et al. (2011a); Fert et al. (2013), spin-orbit torque devices Miron et al. (2011b); Garello et al. (2013), spin filters and pumps Koga et al. (2002), dissipation-less spin transport Sinova et al. (2015) and interaction-based quantum transport Kumar and Pepper (2021). In addition, Rashba SOC plays a significant role in emerging quantum technologies from solid state quantum simulators Reyren et al. (2007); Caviglia et al. (2010); Brinkman et al. (2007); Ben Shalom et al. (2010) to fault-tolerant quantum computation through formation of Majorana fermions Alicea (2012) and information encoding in pseudo-spins Xu et al. (2014). Therefore, precise knowledge of the SOC parameters is necessary for studying both equilibrium and non-equilibrium properties of the system. So far, SOC has been measured using various methods such as, electron transport Knap et al. (1996); Luo et al. (1990), spectroscopy LaShell et al. (1996); Nitta et al. (1997), and spin Hall effect Wunderlich et al. (2005). The fundamental and technological importance of SOC and the recent developments in the field of quantum sensing pose a timely question: can the precision, with which the SOC parameters are measured, be enhanced using quantum features?
@@
Quantum sensing has now established itself as a key component of quantum technologies that can surpass the limits of traditional classical sensors Degen et al. (2017). The Cramér-Rao inequality from estimation theory provides a lower bound on the precision of estimating an unknown parameter in terms of inverse of the Fisher information. Generally, as the probe size $L$ is increased, Fisher information grows as $F{\sim}L^{\beta}$. For classical probes, the best achievable scaling is the standard limit where $\beta{=}1$. Meanwhile, exploiting quantum features can improve the precision beyond the capacity of classical probes, with $\beta{>}1$ Paris (2009); Degen et al. (2017); Braun et al. (2018), which is known as quantum-enhanced sensitivity. The concept of quantum-enhanced sensitivity was first introduced in an interferometric setup utilizing Greenberger–Horne–Zeilinger (GHZ) type entangled states Giovannetti et al. (2004, 2006, 2011). It significantly improves the precision of phase shift measurements to achieve the Heisenberg limit, $\beta{=}2$. Since then, this phenomenon has been verified in various platforms including ion-traps Leibfried et al. (2004), optical systems Mitchell et al. (2004); Pezzé et al. (2007); Ono et al. (2013), superconducting qubits Wang et al. (2019) and nitrogen-vacancy centers Bonato et al. (2016). However, the approach exclusively relies on GHZ-type entangled states, which are challenging to generate and highly susceptible to decoherence and particle loss. Aside from entanglement, squeezing is another established resource for interferometric phase estimation, especially with photonic setups Pezzé and Smerzi (2008); Schnabel (2017); Polino et al. (2020) and spin squeezed systems Ma and Wang (2009); Frérot and Roscilde (2018); Ma et al. (2011), see Ref. Maccone and Riccardi (2020) for a unified framework of squeezing-based metrology. These methods perform with a phase generating operation to encode information into the quantum state of the probe. Any disturbance in this unitary process reduces the precision, therefore limits the interferometric sensing method to scale up effectively De Pasquale et al. (2013).
@@
Another approach to quantum sensing utilizes many-body probes by leveraging a variety of quantum features. Although interactions have destructive effects in the GHZ-based sensing schemes, they play the primary role in many-body sensors Montenegro et al. (2025). In particular, quantum criticality has been identified as a resource for achieving quantum enhanced sensitivity. This includes various types of criticalities, such as first-order Raghunandan et al. (2018); Mirkhalaf et al. (2020); Yang and Jacob (2019); Heugel et al. (2019); Sarkar et al. (2025), second-order Zanardi and Paunković (2006); Zanardi et al. (2007, 2008); Invernizzi et al. (2008); Gu (2010); Gammelmark and Mølmer (2011); Skotiniotis et al. (2015); Rams et al. (2018); Chen et al. (2021); Chu et al. (2021); Liu et al. (2021); Montenegro et al. (2021), Floquet Mishra and Bayat (2021, 2022), time crystal Montenegro et al. (2023); Iemini et al. (2024); Yousefjani et al. (2025a); Gribben et al. (2025); Shukla et al. (2025), Stark He et al. (2023); Yousefjani et al. (2025b, 2023) and quasi-periodic Sahoo et al. (2024) localization, and topological Sarkar et al. (2022, 2024); Mukhopadhyay and Bayat (2024) phase transitions. In all these criticalities, a common theme that emerges as the resource for quantum enhancement of the sensing capability is the closing of the energy gap at the phase transition point Montenegro et al. (2025). Nonetheless, the phase transition occurs only at a particular value of a Hamiltonian parameter which restricts the quantum advantage to a small region around the critical point. As one moves away from the criticality, the energy gap reopens and the quantum enhancement starts to disappear, typically falling back to the standard limit of sensing. This necessitates fine-tuning, making it primarily useful for local sensing where substantial prior knowledge about the parameter of interest is required. Therefore, it is desirable to have a quantum system that feature the gap closing behavior over a wide range of the parameter to be estimated.
@@
II Overview: parameter estimation
II.1 Cramér-Rao bound
To estimate $d$ number of unknown parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})$, the first step is to encode them in a quantum probe whose density operator $\rho_{\bm{\alpha}}$ depends on $\bm{\alpha}$. To determine these parameters, one needs to perform a measurement, which is represented by a set of Positive Operator-Valued Measure (POVM) operators $\{\Pi_{k}\}$ applied to the probe. The outcomes follow a classical probability distribution, where the probability of the $k$-th outcome is given by $p_{k}(\bm{\alpha})=\mathrm{Tr}[\Pi_{k}\rho_{\bm{\alpha}}]$. These outcomes are fed into an estimator function to infer the parameters. For this measurement setup, the uncertainty in estimating $\bm{\alpha}$ can be quantified by the covariance matrix $\text{Cov}({\bm{\alpha}})]_{i,j}{=}\braket{\alpha_{i}\alpha_{j}}-\braket{\alpha_{i}}\braket{\alpha_{j}}$ . Note that the diagonal elements are the single parameter variances $\sigma^{2}_{\alpha_{i}}$. The lower bound in uncertainty is set by the Cramér-Rao inequality Rao (1992); Braunstein and Caves (1994); Cramér (1999); Liu et al. (2019)
@@ II.1 Cramér-Rao bound
where $M$ is the number of measurements and $\mathcal{F}^{C}$ is the $d\times d$ classical Fisher information matrix (CFIM) with elements
@@
This basis-dependent lower bound can be optimized over all possible measurements to give the ultimate obtainable precision in terms of the quantum Cramér-Rao bound Liu et al. (2019)
@@
The quantum Fisher information matrix (QFIM) elements can be expressed with the eigen-decomposition of the density operator $|\rho_{\bm{\alpha}}{=}\sum_{n,(\lambda_{n}\neq 0)}\lambda_{n}\ket{\lambda_{n}}\bra{\lambda_{n}}$ in the following way Liu et al. (2019),
@@
$+\sum_{n,(\lambda_{n}\neq 0)}4\lambda_{n}\text{Re}(\braket{\partial_{\alpha_{i}}\lambda_{n}|\partial_{\alpha_{j}}\lambda_{n}})$
@@
$-\sum_{\begin{subarray}{c}n,m\\ (\lambda_{n},\lambda_{m}\neq 0)\end{subarray}}8\frac{\lambda_{n}\lambda_{m}}{\lambda_{n}+\lambda_{m}}\text{Re}(\braket{\partial_{\alpha_{i}}\lambda_{n}|\lambda_{m}}\braket{\lambda_{m}|\partial_{\alpha_{j}}\lambda_{n}}).$ (4)
For a pure state $|\rho_{\bm{\alpha}}=\ket{\psi_{\bm{\alpha}}}\bra{\psi_{\bm{\alpha}}}$, QFIM becomes
$\mathcal{F}^{Q}_{i,j}=4\text{Re}(\braket{\partial_{\alpha_{i}}\psi|\partial_{\alpha_{j}}\psi})-\braket{\partial_{\alpha_{i}}\psi|\psi}\braket{\psi|\partial_{\alpha_{j}}\psi}).$ (5)
@@
To extract a scalar form from the matrix inequalities, a positive weight matrix $W$ can be multiplied on both sides, followed by a trace operation. In particular, for the choice of $W$ being an identity matrix one gets $\sum_{i}\sigma^{2}_{\alpha_{i}}\geq\text{Tr}((\mathcal{F}^{Q})^{-1})/M$. Another useful choice is to take a $W$ matrix with all elements being zero except $W_{i,i}{=}1$. Then the Cramér-Rao inequality becomes $\sigma^{2}_{\alpha_{i}}\geq(\mathcal{F}^{Q})^{-1}_{i,i}/M\geq 1/M\mathcal{F}^{Q}_{i,i}$, in which the last term is associated with the uncertainty in $\alpha_{i}$ when all the other parameters are known. This situation is known as the single parameter estimation problem where the error in estimating one unknown parameter $\alpha$ is lower bounded by $\sigma^{2}_{\alpha}\geq 1/MF^{C}(\alpha)\geq 1/MF^{Q}(\alpha)$. Here, $F^{C}(\alpha)$ and $F^{Q}(\alpha)$ are the single parameter classical Fisher information (CFI) and quantum Fisher information (QFI), and are just the diagonal elements in CFIM and QFIM, respectively. One choice of optimal basis, in which CFI saturates the QFI, is given by the projectors to the eigenstates of the so called symmetric logarithmic derivative operator (SLD) $\mathcal{L}$, which are implicitly defined for a probe state $\rho$ as $\partial_{\alpha}\rho{=}(\rho\mathcal{L}{+}\mathcal{L}\rho)/2$.
II.2 Bayesian estimation
The above description, known as the frequentist approach, expresses the lower bound of the precision in the case of ideal estimation. A more practical and measurement data based approach is given by the Bayesian analysis. In this approach, the parameters to be estimated ($\bm{\alpha}$) are treated as random variables. An initial knowledge of the parameters is required in terms of a probability distribution, which is given by the prior $P(\bm{\alpha})$. If no prior information about the parameters are known except for their ranges, i.e. $\alpha_{i}\in\left[\alpha_{i,\rm{min}},\alpha_{i,\rm{max}}\right]$, then the prior is a flat distribution $P(\bm{\alpha})=\prod_{i}\frac{1}{\alpha_{i,\rm{max}}-\alpha_{i,\rm{min}}}$. The experimental data is acquired by repeating the measurement process $M$ times where the $k$-th outcome is obtained $n_{k}$ times, with $\sum_{k}n_{k}{=}M$. From this data, one can construct the likelihood which is the conditional probability distribution $P(\{n_{k}\}|\bm{\alpha})$ for getting a dataset $\{n_{k}\}$ given that the parameters were $\bm{\alpha}$. This is given by the multinomial distribution, $P(\{n_{k}\}|\bm{\alpha})=\frac{M!}{\prod_{k}n_{k}!}\prod_{k}p_{k}^{n_{k}}$, where $p_{k}$ is the theoretical probability of the $k$-th outcome. The likelihood function compares the theoretical probability distribution $\{p_{k}\}$ with the experimentally obtained distribution $\{n_{k}/M\}$. Now Bayes’ theorem can be applied to write down the posterior which is the conditional probability distribution for the parameters having values $\bm{\alpha}$ given the experimental outcomes,
$P(\bm{\alpha}|\{n_{k}\})=\frac{P(\{n_{k}\}|\bm{\alpha})\,P(\bm{\alpha})}{\mathcal{N}},$ (6)
where $\mathcal{N}$ is the normalization factor and is the probability of obtaining the experimental outcome. This does not depend on $\bm{\alpha}$ and hence does not affect further analysis. In the limit of large $M$, the central limit theorem applies, and the posterior takes a Gaussian form, the mean and variance of which give the estimated parameter values and precisions, respectively.
II.3 Signal-to-noise ratio
A popular and experimentally relevant tool to express the sensitivity of a sensing procedure is given by the signal-to-noise ratio (SNR), which can be written as $\alpha/\sigma_{\alpha}$ in a single parameter scenario Garbe et al. (2020). This ratio of the parameter to be estimated and the uncertainty, given by the precision achieved, quantifies the sensitivity. In the optimal scenario, the Cramér-Rao bound can be saturated for large $M$ and thus SNR $\approx\alpha\sqrt{MF^{Q}(\alpha)}$. The dependence on both the parameter value and the error makes SNR convey additional information than QFI. Especially, in the context of weak field sensing, where a much smaller error is required, the SNR value becomes particularly important. In the multi-parameter case, this can be generalized Mihailescu et al. (2024) to express the sensitivity for the $\alpha_{i}$ parameter as $\alpha_{i}\sqrt{M/(\mathcal{F}^{Q})^{-1}_{i,i}}$. In this work, we showcase both these scenarios with the aid of this optimal SNR, as well as with the Bayesian case where the estimated value and precision is determined from measurement data.
II.4 Numerical details
[…]
III Model: 1D quantum wire
@@
The model considered in this work is based on the transport of electrons in a one-dimensional ballistic quantum wire in the presence of Rashba SOC Středa and Šeba (2003); Birkholz and Meden (2008). Low dimensional transport properties of electrons are of extreme importance in condensed matter physics, with applications in spintronics and quantum information processing. Usually, such SOC is associated with two-dimensional electron gas systems in xyx-y plane with strong confinement along the $z$-axis. However, it is also prevalent to investigate interesting effects arising from motions in even reduced dimensions. Further confinement to a wire geometry can enhance the effect of SOC Datta and Das (1990) and has led to extensive studies on the transports of non-interacting electrons in quasi 1D systems. The effective dynamics can be described by a tight-binding lattice model, which is governed by the Hamiltonian in the following form
@@
$+\alpha_{y}\sum_{j,\sigma,\sigma^{\prime}}(i\sigma_{z})_{\sigma,\sigma^{\prime}}c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma^{\prime}}+\text{H.c.},$ (9)
@@
IV Single Parameter Sensing
@@
In low-dimensional transport experiments, accurate knowledge of the SOC parameters is desired. Therefore, we first focus on estimating the SOC parameter in the uniform Rashba field scenario along $y$- and $z$-direction, i.e. $\alpha_{y}{=}\alpha_{z}$. The non-uniform SOC case will be addressed in the next section. The goal is to analyze how well one can estimate $\alpha_{z}$ using the ground state of the system as a probe. To stay close to the experimental systems, we consider finite wires with $L$ lattice sites and study the growth of QFI with system size. As the scaling of QFI is closely connected to the gap closing of the system, we also analyze the energy gap $\Delta$ between first excited state and ground state. In our system, we also notice that the ground state is doubly degenerate, so a Zeeman term is essential to break the degeneracy. This term need to be strong enough to capture the actual energy gap and not the Zeeman splitting. For example, one should choose B>0.005tB{>}0.005t for L=100L{=}100.
IV.1 Single particle probe